find the slope of

(1/3,-3/5) (-5/6,7/10)

Answer by lwsshak3(11628)

find the slope of
(1/3,-3/5) (-5/6,7/10)
slope=∆y/∆x=(7/10-(-3/5))/(-5/6-1/3)
=(7/10+6/10))/(-5/6-2/6)
=(13/10)/(-7/6)
=-78/70
=-39/35